### v2

@@ 5.2 K-means experiments
To ensure accurate performance measurement, a 5-fold cross-validation is performed with a fixed seed for each experiment. The number of clusters is set to i.e. the number of classes in Table 2. K-means are first initialized with the k-means++ algorithm Arthur & Vassilvitskii (2007), based on a sampling of the empirical probability distributions of the points (of the entire dataset) that accelerates convergence. All K-means models (issued from the 12 distances) are then initialized with the same optimized centroids. For Gini K-means, a grid search is performed over the hyper-parameter , selecting the value that maximized the precision score. Indeed, since the UCI datasets are supervised, computing the silhouette criterion is not required. After predicting the labels on the test data, the Kuhn-Munkres algorithm Kuhn (1955) is applied to align the predicted clusters with the true labels, as K-means does not assign labels corresponding to the actual classes. Finally, the average precision and recall are computed across all folds.
Without Noise. The generalized Gini prametric achieves the best performance thanks to its hyper-parameter. It is top-ranked on 5 datasets out of 16 (in terms of precision Table 26), with a mean rank of 3.5 (the best mean rank of 4.19 for recall Table 27). The second best method is the K-means based on the Hassanat distance in terms of precision and recall (mean rank of 5.06 and 4.75 respectively). The details about precision and recall are relegated in Tables 7-11 of Appendix B.
@@
With 5% and 10% Noise. The same experiments are conducted for different levels of noise (5% and 10%) as in the KNN experiments, under the same conditions. The Gini K-means reaches the top position 3 times (out of 16 datasets) with a mean ranking of 3.75 (5% noise). The Hellinger distance is ranked 2nd, with more top-ranked positions (5 out of 16 datasets), but with a larger variability so that the mean rank is only 5 (like Hassanat). In the case of 10% noise, the generalized Gini prametric has only 2 top-ranked positions and a mean rank of 4.19. Again, Hellinger has much more top-ranked positions (5 out of 16) but a mean rank of 5.06, exhibiting a more important variability in the ranking (sometimes ranked 10 or 11). The results are presented in Appendix B (Tables 12-20).
Strategy to select for real applications. The performances described above have been computed with known labels for the test points. In real applications, once is fixed, a strategy is necessary to select the hyper-parameter. As described in Algorithm 1, a 5-fold cross-validation with a fixed seed is used to select the hyper-parameter that maximizes the silhouette score (mean across all folds). Tables 38 and 39 indicate that the generalized Gini prametric is ranked 4th in precision and 3rd in recall (for ). This experiment shows that maximizing the silhouette score is not the best method, but it allows the generalized Gini prametric to be ranked among the top 4 distances (Tables 21-24 in Appendix B).
@@ 5.3 Illustration on Agglomerative Hierarchical Clustering
For this additional experiment, noise has been directly added to the data using a standard normal distribution with a noise level of 10% to evaluate which distance metric demonstrates greater robustness under noisy conditions. A 3-fold cross-validation has been made and the linkage criterion has been fixed to ‘average’, i.e., the average distance between all points between clusters is used for all methods. The generalized Gini prametric is top-ranked with 7 top positions out of 16 datasets, and a mean rank of 1.94 (for precision). However, for the recall Ward’s technique is the best (7 top positions and an average rank of 2.62), see Appendix A.5.
6 Conclusion
The generalized Gini prametric, while not being a true distance, satisfies two invariance properties, making it robust to outliers. In particular, it is linearly invariant (similar to the Euclidean distance) and rank invariant. Accordingly, monotonic increasing transformations preserve ranks, ensuring that the Gini prametric exhibits small variations. Consequently, when the data contain outliers or measurement errors, the generalized Gini prametric helps achieve robust results, as shown on the 16 UCI datasets.
The experiments on the UCI datasets highlight the limitations of the generalized Gini prametric in unsupervised tasks. In particular, a strategy for selecting the hyper-parameter, such as using the silhouette score, must be chosen without prior knowledge of the quality of the results. Further grid searches for the hyper-parameter remain an open problem.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning with the Gini methodology. This method can help to refine standard clustering approach, particularly in the presence of outliers.
Appendix A Appendix
A.1 Proposition 4.1
The proof is exactly the same as the one of Cover (1967). The convergence is true for any type of metric used and the bound on the error rate does not change.
A.2 Proposition 4.2
The last inequality comes from the algorithm itself. Indeed, if , this is because . Now, we must prove that taking the sum over on the last expression implies:
with . Taking the derivative of with respect to , and noting that for small variations of ranks remain constant, yields
Using the scalar product this is equivalent to:
Thus,
And finally,
A.3 Gini K-means without noise
| Gini Generalized | 1.18 | 0.4915 | 0.4161 | 4.88 | 0.6339 | 0.6363 | 2.48 | 0.6252 | 0.5650 | 1.72 | 0.6186 | 0.3495 | 1.84 | 0.8214 | 0.5520 | 2.14 | 0.4772 | 0.3409 | 2.36 | 0.4466 | 0.1273 | 1.62 | 0.7052 | 0.0759 |
| Gini | - | 0.3897 | 0.3535 | - | 0.5822 | 0.5816 | - | 0.4645 | 0.5360 | - | 0.4916 | 0.3330 | - | 0.7131 | 0.5408 | - | 0.4082 | 0.3475 | - | 0.1467 | 0.1273 | - | 0.6828 | 0.1174 |
| Euclidean | - | 0.3147 | 0.2882 | - | 0.6077 | 0.6115 | - | 0.4123 | 0.5130 | - | 0.4239 | 0.2672 | - | 0.6977 | 0.5296 | - | 0.3183 | 0.3213 | - | 0.5445 | 0.1227 | - | 0.4288 | 0.0155 |
| Manhattan | - | 0.3586 | 0.3280 | - | 0.6083 | 0.6123 | - | 0.4123 | 0.5130 | - | 0.4113 | 0.2573 | - | 0.7204 | 0.5728 | - | 0.3247 | 0.3443 | - | 0.5992 | 0.0977 | - | 0.4288 | 0.0155 |
| Minkowski | 3.0 | 0.3200 | 0.2910 | 3.0 | 0.6085 | 0.6123 | 3.0 | 0.4123 | 0.5130 | 3.0 | 0.4254 | 0.2605 | 3.0 | 0.6475 | 0.4928 | 3.0 | 0.3208 | 0.2909 | 3.0 | 0.5315 | 0.1386 | 3.0 | 0.4288 | 0.0155 |
| Cosine | - | 0.4013 | 0.3621 | - | 0.5974 | 0.6006 | - | 0.3084 | 0.4840 | - | 0.5479 | 0.2934 | - | 0.7292 | 0.5088 | - | 0.3259 | 0.3770 | - | 0.5952 | 0.4886 | - | 0.6376 | 0.1575 |
| Canberra | - | 0.5419 | 0.4505 | - | 0.5901 | 0.5948 | - | 0.5435 | 0.5250 | - | 0.5651 | 0.2013 | - | 0.4913 | 0.3744 | - | 0.6109 | 0.4388 | - | 0.2348 | 0.1136 | - | 0.6864 | 0.1798 |
| Hellinger | - | 0.5754 | 0.5021 | - | 0.6015 | 0.5868 | - | 0.3364 | 0.4740 | - | 0.4159 | 0.1985 | - | 0.5745 | 0.3760 | - | 0.4612 | 0.4304 | - | 0.3319 | 0.2273 | - | 0.7150 | 0.0414 |
| Jensen-Shannon | - | 0.1289 | 0.3590 | - | 0.3085 | 0.5554 | - | 0.3095 | 0.5050 | - | 0.2548 | 0.0363 | - | 0.4687 | 0.1808 | - | 0.2361 | 0.0230 | - | 0.2621 | 0.1523 | - | 0.7150 | 0.0588 |
| PearsonChi2 | - | 0.1289 | 0.3590 | - | 0.5043 | 0.5094 | - | 0.2789 | 0.4960 | - | 0.4802 | 0.2644 | - | 0.4578 | 0.3616 | - | 0.5301 | 0.4550 | - | 0.5397 | 0.4955 | - | 0.6623 | 0.1678 |
| VicisSymmetric1 | - | 0.5412 | 0.5070 | - | 0.5225 | 0.5109 | - | 0.2954 | 0.4970 | - | 0.3202 | 0.1551 | - | 0.4240 | 0.3840 | - | 0.5324 | 0.4189 | - | 0.3484 | 0.1250 | - | 0.5064 | 0.4039 |
| Hassanat | - | 0.4226 | 0.3821 | - | 0.5556 | 0.5517 | - | 0.5754 | 0.5750 | - | 0.5199 | 0.4033 | - | 0.5938 | 0.4464 | - | 0.5128 | 0.4769 | - | 0.3946 | 0.3250 | - | 0.6958 | 0.1400 |
| Gini Generalized | 3 3.52 | 0.0998 | 0.1 | 3 4.10 | 0.1913 | 0.1738 | 2 2.70 | 0.9234 | 0.9157 | 2 1.22 | 0.6339 | 0.5956 | 3 3.12 | 0.3741 | 0.2978 | 3 4.52 | 0.2946 | 0.2801 | 2 1.66 | 0.7693 | 0.6145 | 6 2.54 | 0.1492 | 0.0699 |
| Gini | 3 2 | 0.0828 | 0.1 | 3 2 | 0.1172 | 0.1005 | 2 2 | 0.9129 | 0.9052 | 2 2 | 0.5985 | 0.5618 | 3 2 | 0.3419 | 0.2644 | 3 2 | 0.2596 | 0.2517 | 2 2 | 0.6891 | 0.6101 | 6 2 | 0.0584 | 0.0327 |
| Euclidean | 3 - | 0.0757 | 0.0933 | 3 - | 0.0610 | 0.0502 | 2 - | 0.8802 | 0.8524 | 2 - | 0.5636 | 0.5373 | 3 - | 0.3520 | 0.2349 | 3 - | 0.2088 | 0.2269 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0388 | 0.0186 |
| Manhattan | 3 - | 0.0788 | 0.1 | 3 - | 0.0680 | 0.0613 | 2 - | 0.8782 | 0.8489 | 2 - | 0.5818 | 0.5518 | 3 - | 0.3531 | 0.2567 | 3 - | 0.2057 | 0.2234 | 2 - | 0.4873 | 0.5590 | 6 - | 0.0628 | 0.0279 |
| Minkowski | 3 3 | 0.0774 | 0.0933 | 3 3 | 0.0643 | 0.0502 | 2 3 | 0.8812 | 0.8542 | 2 3 | 0.5784 | 0.5419 | 3 3 | 0.3498 | 0.2311 | 3 3 | 0.2100 | 0.2281 | 2 3 | 0.4873 | 0.5590 | 6 3 | 0.0359 | 0.0186 |
| Cosine | 3 - | 0.0188 | 0.02 | 3 - | 0.1370 | 0.1119 | 2 - | 0.9128 | 0.9051 | 2 - | 0.5627 | 0.5325 | 3 - | 0.3945 | 0.3055 | 3 - | 0.1984 | 0.2186 | 2 - | 0.6770 | 0.6725 | 6 - | 0.0599 | 0.0280 |
| Canberra | 3 - | 0.0386 | 0.0467 | 3 - | 0.3990 | 0.3430 | 2 - | 0.9237 | 0.9227 | 2 - | 0.5407 | 0.5225 | 3 - | 0.4449 | 0.3979 | 3 - | 0.2292 | 0.2234 | 2 - | 0.5208 | 0.5203 | 6 - | 0.1123 | 0.0421 |
| Hellinger | 3 - | 0.2906 | 0.2933 | 3 - | 0.2646 | 0.2740 | 2 - | 0.7313 | 0.7156 | 2 - | 0.3424 | 0.4901 | 3 - | 0.5476 | 0.3854 | 3 - | 0.3701 | 0.3275 | 2 - | 0.6701 | 0.5725 | 6 - | 0.1579 | 0.1728 |
| JensenShannon | 3 - | 0.4642 | 0.46 | 3 - | 0.0543 | 0.1352 | 2 - | 0.1272 | 0.3268 | 2 - | 0.5549 | 0.5531 | 3 - | 0.5241 | 0.5378 | 3 - | 0.3695 | 0.4196 | 2 - | 0.6830 | 0.6319 | 6 - | 0.0 | 0.0 |
| PearsonChi2 | 3 - | 0.2688 | 0.2867 | 3 - | 0.4505 | 0.4865 | 2 - | 0.8890 | 0.8683 | 2 - | 0.5973 | 0.5823 | 3 - | 0.3998 | 0.2990 | 3 - | 0.3644 | 0.3509 | 2 - | 0.6705 | 0.5913 | 6 - | 0.2025 | 0.1457 |
| VicisSymmetric1 | 3 - | 0.0667 | 0.0333 | 3 - | 0.2341 | 0.2037 | 2 - | 0.8656 | 0.8420 | 2 - | 0.4488 | 0.5296 | 3 - | 0.4766 | 0.2786 | 3 - | 0.4819 | 0.4031 | 2 - | 0.6502 | 0.5652 | 6 - | 0.1515 | 0.1260 |
| Hassanat | 3 - | 0.0497 | 0.06 | 3 - | 0.3990 | 0.3598 | 2 - | 0.9240 | 0.9157 | 2 - | 0.5834 | 0.5470 | 3 - | 0.3772 | 0.2490 | 3 - | 0.3563 | 0.2931 | 2 - | 0.7858 | 0.7739 | 6 - | 0.0553 | 0.0279 |
A.4 Gini KNN without noise
| Euclidean | 2 | 0.89 | 0.89 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 1 | 0.56 | 0.57 | 10 | 0.84 | 0.88 | 8 | 0.75 | 0.75 | 9 | 0.91 | 0.90 | 2 | 0.60 | 0.37 |
| Manhattan | 1 | 0.91 | 0.90 | 1 | 1.00 | 1.00 | 2 | 0.70 | 0.66 | 4 | 0.56 | 0.58 | 8 | 0.86 | 0.87 | 10 | 0.74 | 0.76 | 9 | 0.91 | 0.91 | 2 | 0.60 | 0.37 |
| Minkowski | 2 | 0.88 | 0.87 | 2 | 1.00 | 1.00 | 2 | 0.68 | 0.65 | 3 | 0.56 | 0.50 | 10 | 0.85 | 0.88 | 8 | 0.76 | 0.76 | 9 | 0.90 | 0.89 | 2 | 0.59 | 0.36 |
| Cosine | 2 | 0.89 | 0.89 | 1 | 0.99 | 0.99 | 8 | 0.69 | 0.69 | 6 | 0.52 | 0.55 | 3 | 0.93 | 0.90 | 2 | 0.74 | 0.76 | 10 | 0.88 | 0.88 | 1 | 0.55 | 0.36 |
| Hassanat | 2 | 0.92 | 0.92 | 1 | 1.00 | 1.00 | 10 | 0.73 | 0.71 | 3 | 0.67 | 0.67 | 3 | 0.84 | 0.82 | 4 | 0.74 | 0.76 | 10 | 0.91 | 0.90 | 2 | 0.58 | 0.37 |
| Canberra | 6 | 0.92 | 0.91 | 1 | 0.93 | 0.93 | 3 | 0.70 | 0.70 | 9 | 0.66 | 0.71 | 3 | 0.84 | 0.79 | 5 | 0.73 | 0.75 | 6 | 0.92 | 0.90 | 2 | 0.58 | 0.36 |
| Hellinger | 2 | 0.90 | 0.90 | 1 | 0.95 | 0.95 | 8 | 0.70 | 0.69 | 8 | 0.65 | 0.69 | 8 | 0.87 | 0.88 | 7 | 0.73 | 0.76 | 5 | 0.93 | 0.93 | 2 | 0.60 | 0.36 |
| Jensen-Shannon | 2 | 0.78 | 0.40 | 1 | 0.89 | 0.89 | 10 | 0.71 | 0.70 | 10 | 0.63 | 0.67 | 1 | 0.94 | 0.94 | 7 | 0.74 | 0.75 | 9 | 0.89 | 0.88 | 2 | 0.72 | 0.06 |
| PearsonChi2 | 1 | 0.41 | 0.64 | 1 | 0.95 | 0.95 | 8 | 0.73 | 0.71 | 7 | 0.63 | 0.67 | 8 | 0.87 | 0.89 | 8 | 0.72 | 0.75 | 7 | 0.92 | 0.92 | 2 | 0.58 | 0.36 |
| VicisSymmetric1 | 7 | 0.81 | 0.76 | 1 | 0.92 | 0.92 | 8 | 0.70 | 0.70 | 10 | 0.62 | 0.67 | 8 | 0.85 | 0.88 | 5 | 0.74 | 0.71 | 8 | 0.91 | 0.91 | 2 | 0.60 | 0.35 |
| Gini | 2 2 | 0.85 | 0.84 | 1 2 | 1.00 | 1.00 | 6 2 | 0.71 | 0.70 | 1 2 | 0.61 | 0.61 | 10 2 | 0.85 | 0.89 | 5 2 | 0.71 | 0.69 | 5 | 0.92 | 0.92 | 2 | 0.58 | 0.37 |
| Generalized Gini | 2 4.89 | 0.90 | 0.89 | 1 1.96 | 1.00 | 1.00 | 7 3.35 | 0.73 | 0.73 | 1 2.12 | 0.63 | 0.63 | 7 5.09 | 0.87 | 0.89 | 5 5.09 | 0.72 | 0.75 | 8 5.09 | 0.93 | 0.93 | 2 2.67 | 0.58 | 0.37 |
| Euclidean | 5 - | 0.97 | 0.97 | 1 - | 0.71 | 0.71 | 10 - | 0.93 | 0.93 | 1 - | 0.84 | 0.83 | 6 - | 0.64 | 0.65 | 1 - | 0.87 | 0.87 | 6 - | 0.70 | 0.69 | 3 - | 0.74 | 0.72 |
| Manhattan | 1 - | 0.96 | 0.96 | 1 - | 0.77 | 0.77 | 6 - | 0.94 | 0.94 | 1 - | 0.86 | 0.86 | 9 - | 0.64 | 0.65 | 1 - | 0.90 | 0.90 | 8 - | 0.75 | 0.73 | 1 - | 0.75 | 0.71 |
| Minkowski | 5 - | 0.97 | 0.97 | 1 - | 0.68 | 0.69 | 10 - | 0.93 | 0.93 | 3 - | 0.84 | 0.83 | 8 - | 0.63 | 0.64 | 1 - | 0.85 | 0.85 | 10 - | 0.70 | 0.67 | 2 - | 0.68 | 0.65 |
| Cosine | 7 - | 0.98 | 0.97 | 1 - | 0.84 | 0.84 | 7 - | 0.93 | 0.92 | 1 - | 0.86 | 0.85 | 9 - | 0.63 | 0.63 | 1 - | 0.89 | 0.88 | 9 - | 0.72 | 0.72 | 8 - | 0.69 | 0.69 |
| Hassanat | 10 - | 0.95 | 0.95 | 1 - | 0.96 | 0.95 | 7 - | 0.96 | 0.96 | 1 - | 0.87 | 0.87 | 8 - | 0.63 | 0.64 | 7 - | 0.89 | 0.89 | 8 - | 0.86 | 0.84 | 4 - | 0.74 | 0.68 |
| Canberra | 9 - | 0.95 | 0.95 | 1 - | 0.97 | 0.97 | 6 - | 0.96 | 0.95 | 1 - | 0.85 | 0.83 | 7 - | 0.65 | 0.65 | 1 - | 0.85 | 0.85 | 6 - | 0.87 | 0.86 | 2 - | 0.67 | 0.68 |
| Hellinger | 2 - | 0.97 | 0.97 | 4 - | 0.91 | 0.92 | 4 - | 0.95 | 0.95 | 1 - | 0.87 | 0.86 | 6 - | 0.64 | 0.65 | 1 - | 0.86 | 0.86 | 2 - | 0.77 | 0.76 | 1 - | 0.69 | 0.69 |
| Gini | 4 2 | 0.97 | 0.97 | 5 2 | 0.90 | 0.89 | 5 2 | 0.95 | 0.95 | 1 2 | 0.87 | 0.87 | 7 - | 0.66 | 0.67 | 1 2 | 0.93 | 0.93 | 6 2 | 0.77 | 0.75 | 3 2 | 0.73 | 0.71 |
| Generalized Gini | 7 1.1 | 0.98 | 0.98 | 3 4.45 | 0.91 | 0.91 | 4 3.11 | 0.95 | 0.95 | 1 2.6 | 0.88 | 0.88 | 7 2.39 | 0.67 | 0.67 | 1 3.48 | 0.93 | 0.93 | 5 3.45 | 0.79 | 0.79 | 3 1.37 | 0.77 | 0.74 |
| JensenShannon | 3 - | 0.98 | 0.96 | 10 - | 0.95 | 0.94 | 3 - | 0.93 | 0.93 | 1 - | 0.87 | 0.86 | 9 - | 0.64 | 0.65 | 9 - | 0.86 | 0.86 | 9 - | 0.77 | 0.78 | 1 - | 0.67 | 0.66 |
| PearsonChi2 | 2 - | 0.97 | 0.97 | 4 - | 0.92 | 0.92 | 4 - | 0.95 | 0.95 | 1 - | 0.86 | 0.86 | 2 - | 0.57 | 0.59 | 7 - | 0.76 | 0.75 | 7 - | 0.76 | 0.75 | 1 - | 0.68 | 0.67 |
| VicisSymmetric1 | 5 - | 0.97 | 0.95 | 1 - | 0.98 | 0.98 | 5 - | 0.96 | 0.96 | 1 - | 0.79 | 0.78 | 9 - | 0.65 | 0.65 | 5 - | 0.78 | 0.78 | 8 - | 0.81 | 0.79 | 3 - | 0.65 | 0.62 |
A.5 Agglomerative clustering with 10% noise
| Iris | Wine | Breast Cancer | Sonar | Australian | Glass | QSAR | Vehicle | |||||||||||||||||
| Gini Generalized | 3 2.37 | 0.8990 | 0.8867 | 3 4.27 | 0.7864 | 0.7187 | 2 2.50 | 0.9161 | 0.9050 | 2 2.54 | 0.4766 | 0.5336 | 3 2.22 | 0.7647 | 0.5957 | 6 1.20 | 0.5835 | 0.4485 | 3 1.36 | 0.6227 | 0.5982 | 4 4.95 | 0.6631 | 0.6028 |
| Gini | 3 - | 0.8808 | 0.8533 | 3 - | 0.5433 | 0.6575 | 2 - | 0.8697 | 0.8312 | 2 - | 0.3509 | 0.5239 | 3 - | 0.7153 | 0.5826 | 6 - | 0.5573 | 0.4765 | 3 - | 0.3992 | 0.5738 | 4 - | 0.4779 | 0.5035 |
| Euclidean | 3 - | 0.8798 | 0.8133 | 3 - | 0.5077 | 0.6345 | 2 - | 0.7767 | 0.6520 | 2 - | 0.4187 | 0.5239 | 3 - | 0.4434 | 0.5565 | 6 - | 0.4843 | 0.4068 | 3 - | 0.5287 | 0.5982 | 4 - | 0.6847 | 0.5330 |
| Euclidean Ward | 3 - | 0.8773 | 0.8600 | 3 - | 0.7941 | 0.7797 | 2 - | 0.8872 | 0.8632 | 2 - | 0.2947 | 0.5429 | 3 - | 0.7664 | 0.6174 | 6 - | 0.6302 | 0.4648 | 3 - | 0.5360 | 0.5946 | 4 - | 0.6592 | 0.5745 |
| Manhattan | 3 - | 0.8945 | 0.8333 | 3 - | 0.5104 | 0.6511 | 2 - | 0.7767 | 0.6520 | 2 - | 0.3741 | 0.5193 | 3 - | 0.4434 | 0.5565 | 6 - | 0.4024 | 0.4068 | 3 - | 0.5540 | 0.5969 | 4 - | 0.6878 | 0.5366 |
| Minkowski | 3 3 | 0.7308 | 0.6933 | 3 3 | 0.5104 | 0.6511 | 2 3 | 0.7767 | 0.6520 | 2 3 | 0.2848 | 0.5336 | 3 3 | 0.4434 | 0.5565 | 6 3 | 0.4319 | 0.4673 | 3 3 | 0.5657 | 0.6008 | 4 3 | 0.6028 | 0.5485 |
| Cosine | 3 - | 0.6619 | 0.7667 | 3 - | 0.4529 | 0.5901 | 2 - | 0.7831 | 0.6660 | 2 - | 0.2848 | 0.5336 | 3 - | 0.4771 | 0.5580 | 6 - | 0.4272 | 0.4626 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4863 | 0.5047 |
| Canberra | 3 - | 0.4870 | 0.6533 | 3 - | 0.4728 | 0.6518 | 2 - | 0.3940 | 0.6274 | 2 - | 0.4386 | 0.5384 | 3 - | 0.3095 | 0.5551 | 6 - | 0.3332 | 0.4439 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4879 | 0.4941 |
| Hellinger | 3 - | 0.8781 | 0.7667 | 3 - | 0.5793 | 0.6513 | 2 - | 0.7755 | 0.6484 | 2 - | 0.2848 | 0.5336 | 3 - | 0.7562 | 0.5609 | 6 - | 0.4082 | 0.4859 | 3 - | 0.4302 | 0.6033 | 4 - | 0.4774 | 0.5130 |
| JensenShannon | 3 - | 0.4870 | 0.6533 | 3 - | 0.2890 | 0.4878 | 2 - | 0.3940 | 0.6274 | 2 - | 0.2848 | 0.5336 | 3 - | 0.5377 | 0.6101 | 6 - | 0.1265 | 0.3552 | 3 - | 0.3491 | 0.5905 | 4 - | 0.2588 | 0.5071 |
| PearsonChi2 | 3 - | 0.8781 | 0.7667 | 3 - | 0.6765 | 0.6964 | 2 - | 0.6517 | 0.6379 | 2 - | 0.3484 | 0.5143 | 3 - | 0.7562 | 0.5609 | 6 - | 0.4082 | 0.4859 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4757 | 0.5213 |
| VicisSymmetric1 | 3 - | 0.4870 | 0.6533 | 3 - | 0.3309 | 0.4045 | 2 - | 0.3940 | 0.6274 | 2 - | 0.3588 | 0.5193 | 3 - | 0.5318 | 0.6014 | 6 - | 0.4604 | 0.4623 | 3 - | 0.3491 | 0.5905 | 4 - | 0.2984 | 0.5024 |
| Hassanat | 3 - | 0.7904 | 0.7867 | 3 - | 0.6191 | 0.7365 | 2 - | 0.8660 | 0.8314 | 2 - | 0.3458 | 0.5288 | 3 - | 0.6869 | 0.5768 | 6 - | 0.5249 | 0.4907 | 3 - | 0.3491 | 0.5905 | 4 - | 0.4873 | 0.4917 |
@@ A.5 Agglomerative clustering with 10% noise
| Gini Generalized | 7 | 1 | 2 | 2 | 2 | 1 | 1 | 2 | 1 | 2 | 1 | 1 | 2 | 2 | 1 | 3 | 1.94 |
| Gini | 7 | 12 | 3 | 5 | 10 | 4 | 13 | 2 | 3 | 6 | 3 | 6 | 5 | 3 | 7 | 9 | 6.12 |
@@
| Gini Generalized | 7 | 1 | 2 | 3 | 8 | 3 | 1 | 2 | 1 | 3 | 1 | 3 | 4 | 9 | 3 | 1 | 3.25 |
| Gini | 7 | 12 | 4 | 13 | 1 | 3 | 13 | 2 | 3 | 5 | 4 | 9 | 5 | 4 | 13 | 10 | 6.75 |

### v3

@@ 5.2 K-means Experiments
To ensure accurate performance measurement, a 5-fold cross-validation is performed with a fixed seed for each experiment. The number of clusters is set to i.e. the number of classes in Table 2. K-means are first initialized with the k-means++ algorithm (Arthur and Vassilvitskii [4]), based on a sampling of the empirical probability distributions of the points (of the entire dataset) that accelerates convergence. All K-means models (issued from the 12 distances) are then initialized with the same optimized centroids. For Gini K-means, a grid search is performed over the hyper-parameter , selecting the value that maximizes the precision score. Indeed, since the UCI datasets are supervised, computing the silhouette criterion is not required. After predicting the labels on the test data, the Kuhn-Munkres algorithm [24] is applied to align the predicted clusters with the true labels, as K-means does not assign labels corresponding to the actual classes. Finally, the average precision and recall are computed across all folds.
Without Noise. The generalized Gini prametric achieves the best performance thanks to its hyper-parameter. It is top-ranked on 5 datasets out of 16 (in terms of precision Table 6), with a mean rank of 3.5 (the best mean rank of 4.19 for recall Table 7). The second best method is the K-means based on the Hassanat distance in terms of precision and recall (mean rank of 5.06 and 4.75 respectively). The details about precision and recall are relegated in Tables 11-14 of Supplementary Materials [27].
@@
With 5% and 10% Noise. The same experiments are conducted for different levels of noise (5% and 10%) as in the KNN experiments, under the same conditions. The Gini K-means reaches the top position 3 times (out of 16 datasets) with a mean ranking of 3.75 (5% noise). The Hellinger distance is ranked 2nd, with more top-ranked positions (5 out of 16 datasets), but with a larger variability so that the mean rank is only 5 (like Hassanat). In the case of 10% noise, the generalized Gini prametric has only 2 top-ranked positions and a mean rank of 4.19. Again, Hellinger has much more top-ranked positions (5 out of 16) but a mean rank of 5.06, exhibiting a more important variability in the ranking (sometimes ranked 10 or 11). The results are presented in Supplementary Materials (Tables 15-22 [27]).
Strategy to select for real applications. The performances described above have been computed with known labels for the test points. In real applications, once is fixed, a strategy is necessary to select the hyper-parameter. As described in Algorithm 1, a 5-fold cross-validation with a fixed seed is used to select the hyper-parameter that maximizes the silhouette score (mean across all folds). Tables 8 and 9 indicate that the generalized Gini prametric is ranked 4th in precision and 3rd in recall (for ). This experiment shows that maximizing the silhouette score is not the best method, but it allows the generalized Gini prametric to be ranked among the top 4 distances (Tables 23-26 in Supplementary Materials [27]).
@@
Convergence. The performance of Gini K-means is evaluated based on the number of iterations required for convergence. While the grid search used to obtain the optimal parameter () is computationally efficient, it is essential to analyze the convergence time, which is reported in terms of iteration counts. Compared to other methods, the optimized generalized Gini prametric (with ) exhibits an average of 6.5 iterations over the 16 datasets (and over the 5 folds), placing the method among the top five in terms of efficiency.
| Distances | Ir | Wi | Bc | S | A | He | Gl | Ba | G |
| Gini pr. () | 2.4 | 8.8 | 7.6 | 6.6 | 7.4 | 16.6 | 4.8 | 9.2 | 4.8 |
| Gini pr. () | 2.0 | 11.4 | 2.0 | 6.2 | 6.8 | 8.6 | 4.4 | 9.2 | 4.8 |
| Euclidean | 2.2 | 4.4 | 2.8 | 3.4 | 1.8 | 3.0 | 3.2 | 19.4 | 12.2 |
| Manhattan | 2.0 | 7.8 | 2.6 | 4.0 | 1.8 | 3.2 | 3.0 | 12.6 | 12.8 |
| Minkowski | 2.8 | 4.2 | 2.8 | 5.0 | 1.8 | 5.2 | 2.8 | 17.2 | 12.2 |
| Cosine | 1.6 | 1.0 | 1.0 | 2.6 | 1.0 | 1.0 | 2.8 | 3.2 | 1.0 |
| Hassanat | 2.6 | 6.2 | 4.0 | 6.0 | 1.0 | 165.8 | 5.2 | 22.0 | 9.2 |
| Canberra | 3.6 | 6.6 | 3.2 | 8.8 | 300.0 | 21.6 | 9.6 | 83.8 | 12.0 |
| Hellinger | 300.0 | 300.0 | 300.0 | 300.0 | 2.2 | 300.0 | 300.0 | 300.0 | 300.0 |
| Jensen_Shannon | 240.4 | 2.0 | 2.0 | 2.0 | 2.0 | 2.0 | 2.0 | 2.0 | 2.0 |
| Pearson Chi2 | 300.0 | 300.0 | 1.8 | 300.0 | 6.8 | 300.0 | 300.0 | 300.0 | 300.0 |
| Vicis Symmetric | 300.0 | 300.0 | 240.8 | 300.0 | 300.0 | 300.0 | 300.0 | 300.0 | 300.0 |
| Distances | Ha | W | V | B | In | I | QS | Mean |
| Gini prametric () | 3.8 | 18.4 | 12.4 | 8.4 | 7.4 | 4.0 | 5.2 | 8.0 |
| Gini prametric () | 5.0 | 14.2 | 10.8 | 5.4 | 6.0 | 2.4 | 4.6 | 6.5 |
| Euclidean | 5.6 | 4.8 | 3.4 | 2.0 | 2.2 | 2.4 | 3.6 | 4.8 |
| Manhattan | 5.0 | 7.4 | 4.2 | 4.0 | 2.2 | 3.2 | 7.2 | 5.2 |
| Minkowski | 5.2 | 5.0 | 4.4 | 2.0 | 2.2 | 2.2 | 2.4 | 4.8 |
| Cosine | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.4 | 1.0 | 1.4 |
| Hassanat | 19.0 | 1.0 | 18.2 | 6.0 | 4.0 | 5.4 | 11.4 | 17.9 |
| Canberra | 11.4 | 1.0 | 14.4 | 4.8 | 4.4 | 5.0 | 12.8 | 31.4 |
| Hellinger | 300.0 | 3.8 | 300.0 | 300.0 | 3.0 | 300.0 | 300.0 | 244.3 |
| Jensen_Shannon | 2.0 | 62.6 | 2.0 | 2.0 | 2.0 | 2.0 | 2.0 | 20.7 |
| Pearson Chi2 | 243.8 | 257.4 | 300.0 | 264.6 | 86.2 | 2.0 | 300.0 | 222.7 |
| Vicis Symmetric | 300.0 | 2.2 | 300.0 | 300.0 | 300.0 | 300.0 | 300.0 | 277.7 |
@@ 5.3 Illustration on Agglomerative Hierarchical Clustering
For this additional experiment, noise has been directly added to the data using a standard normal distribution with a noise level of 10% to evaluate which distance metric demonstrates greater robustness under noisy conditions. A 3-fold cross-validation has been made and the linkage criterion has been fixed to ‘average’, i.e., the average distance between all points between clusters is used for all methods. The generalized Gini prametric is top-ranked with 7 top positions out of 16 datasets, and a mean rank of 1.94 (for precision).
@@
| Gini prametric | 7 | 1 | 2 | 2 | 2 | 1 | 1 | 2 | 1 | 2 | 1 | 1 | 2 | 2 | 1 | 3 | 1.94 |
| Gini prametric | 7 | 12 | 3 | 5 | 10 | 4 | 13 | 2 | 3 | 6 | 3 | 6 | 5 | 3 | 7 | 9 | 6.12 |
@@
| Gini prametric | 7 | 1 | 2 | 3 | 8 | 3 | 1 | 2 | 1 | 3 | 1 | 3 | 4 | 9 | 3 | 1 | 3.25 |
| Gini prametric | 7 | 12 | 4 | 13 | 1 | 3 | 13 | 2 | 3 | 5 | 4 | 9 | 5 | 4 | 13 | 10 | 6.75 |